(w-12)(aw+9), where a is a real number and a does not equal 0
Answer by jai_kos(139)

(w -12)* (aw+9) = 0
That is (w - 12) = 0 (1)
Or (aw + 9) = 0 (2)

 w = 12

Now substitute w =12 in equation (2), we get
a *12 + 9 = 0
12 a + 9 = 0
12 a = -9
a = -9 /12 = - 3 / 4
Therefore the value for a = (-3/4)